# Generalizable Implicit Motion Modeling for Video Frame Interpolation

### v3

@@ Abstract
Motion modeling is critical in flow-based Video Frame Interpolation (VFI). Existing paradigms either consider linear combinations of bidirectional flows or directly predict bilateral flows for given timestamps without exploring favorable motion priors, thus lacking the capability of effectively modeling spatiotemporal dynamics in real-world videos. To address this limitation, in this study, we introduce Generalizable Implicit Motion Modeling (GIMM), a novel and effective approach to motion modeling for VFI. Specifically, to enable GIMM as an effective motion modeling paradigm, we design a motion encoding pipeline to model spatiotemporal motion latent from bidirectional flows extracted from pre-trained flow estimators, effectively representing input-specific motion priors. Then, we implicitly predict arbitrary-timestep optical flows within two adjacent input frames via an adaptive coordinate-based neural network, with spatiotemporal coordinates and motion latent as inputs. Our GIMM can be smoothly integrated with existing flow-based VFI works without further modifications. We show that GIMM performs better than the current state of the art on the VFI benchmarks.
@@ 1 Introduction
Figure 1: Schematic of motion modeling paradigms in video frame interpolation. (a) A naïvely linear combination of bidirectional flows $F_{0\rightarrow 1},F_{1\rightarrow 0}$ (i.e., flows between input frames) may lead to ambiguous and coarse motion estimation, due to strong overlapped and linear assumptions. (b) A time-condition-based modeling approach may predict suboptimal bilateral flows $F_{t\rightarrow 0},F_{t\rightarrow 1}$ (i.e., flows between estimated and input frames), capturing spatiotemporal changes for moving objects ineffectively. (c) Our generalizable implicit motion modeling can properly represent spatiotemporal dynamics across videos and predict better bilateral flows via an adaptive coordinate-based neural network.
@@
Accurately modeling flow from two distant frames is challenging due to the complexities of capturing subtle and dynamic movements caused by varying motion speeds, object occlusions, and lighting conditions. A commonly used approach in the literature combines bidirectional flows derived from input frames [21, 38, 42, 50, 37] (Figure 1(a)). This method assumes overlapped and linear motion for flow estimation, which does not accurately reflect real-world dynamics. Several studies [19, 28, 52, 40] directly predict bilateral flows based on intermediate discrete timestamps (Figure 1(b)). These approaches model the correlation between frame motions and timestamps without leveraging motion priors within input frames; they thus fall short in capturing complex spatial-temporal changes and handling occluded regions and unexpected deformations. Moreover, this discrete-time-based modeling paradigm is ineffective for arbitrary-time interpolation.
@@
In this study, we explore a more effective and generalizable approach to motion modeling for VFI. Inspired by the success of implicit neural representations [45] in encoding complex high-dimensional data such as 2D images [5], 3D scenes [32], and videos [3], we propose to implicitly model optical flows between two adjacent video frames using coordinate-based neural networks. This implicit paradigm can directly take arbitrary spatiotemporal coordinates as inputs and effectively decode the desired space-time outputs, making it a promising framework for learning highly dynamic optical flows in real-world videos. However, leveraging implicit neural networks for effective motion modeling poses challenges. First, standard implicit neural networks typically perform per-instance modeling, optimizing model parameters for a single specific input. This limitation restricts their applicability across different input video frames. Therefore, we need to develop a more adaptive implicit model capable of capturing motions in any given video. Second, efficiently integrating spatiotemporal information within implicit neural networks is complex, especially when dealing with the intricate motions occurring between two video frames. This necessitates designing appropriate implicit neural architectures that can accurately predict and represent both the spatial and temporal dynamics in videos.
@@
To this end, we propose a novel generalizable implicit flow encoding for motion modeling in VFI, called Generalizable Implicit Motion Modeling (GIMM) (Figure 1(c)). Our method only assumes the availability of bidirectional flows ($F_{0\rightarrow 1},F_{1\rightarrow 0}$ ) of two input frames obtained from a pre-trained optical flow estimator (e.g., RAFT [47], FlowFormer [18]). The input flows can be noisy as this prior will be refined to estimate the bilateral flows ($F_{t\rightarrow 0},F_{t\rightarrow 1}$) between arbitrary intermediate timestamps. Our method is unique in that it can represent input-specific motion priors effectively, for instance, it can accurately capture the complex dynamics of a somersault. This is achieved by introducing a motion encoding pipeline to extract spatiotemporal motion latents from the bidirectional flows. Our method can further enable frame interpolation at arbitrary timestamps, thanks to the adaptive coordinate-based neural network that takes spatiotemporal coordinates and motion latents as inputs. This capability allows our method to generate frames at various temporal granularities, providing flexibility and precision in video frame interpolation.
@@
Our contributions are summarized as follows: We present an effective motion modeling paradigm for video frame interpolation, characterized by a novel generalizable implicit motion modeling framework (GIMM). Our GIMM is capable of accurately predicting optical flow for arbitrary timesteps between two adjacent video frames at any resolution, allowing seamless integration with existing flow-based VFI methods. We demonstrate the advantages of our GIMM in motion modeling for arbitrary-timestep VFI tasks, achieving state-of-the-art performances on various benchmarks.
@@ 2 Related Work
Implicit neural representations. Implicit Neural Representations (INRs) have been shown effective in modeling complex, high-dimensional data for various applications, including video compression [3], novel view synthesis [32, 29], and image super-resolution [5]. Typically, INRs learn a continuous mapping from a set of coordinates to a specific signal using a coordinate-based neural network to implicitly encode data. The flexible and expressive modeling capabilities of INRs motivate us to explore their use to encode intricate motions and capture subtle, dynamic movements of objects in real-world videos. A recent work related to ours is IFE [13]. While both approaches consider implicit flow encoding, our method significantly differs from IFE. Unlike IFE, which focuses on per-scene encoding—where each coordinate-based network is parameterized by a specific video instance—we aim to develop a generalizable motion modeling approach that can be applied across different videos.
@@ 3 Method
This section describes our GIMM framework for video frame interpolation. In Sec. 3.1, we first introduce a typical general flow-based VFI process. We subsequently detail our proposed GIMM framework in Sec. 3.2. We finally present the integration of GIMM for flow-based VFI in Sec. 3.3.
3.1 Preliminary
Video frame interpolation. Given a pair of adjacent video frames $I_{0},I_{1}\in\mathbb{R}^{H\times W\times 3}$ with timesteps $\{0,1\}$, a general flow-based video frame interpolation process is defined as follows:
@@ 3.2 Generalizable Implicit Motion Modeling
Figure 2: Our GIMM first transforms initial bidirectional flows $F_{0\rightarrow 1},F_{1\rightarrow 0}$ as normalized flows $V_{0},V_{1}$. The motion encoder extracts motion features $K_{0},K_{1}$ from $V_{0},V_{1}$, respectively. $K_{0},K_{1}$ are then forward warped at a given timestep $t$ using bidirectional flows to obtain the warped features $K_{t\rightarrow 0},K_{t\rightarrow 1}$. We pass both the warped and initial motion features into a Latent Refiner that outputs motion latent $L_{t}$, representing motion information at $t$. Conditioned on $L_{t}(x,y)$, the coordinate-based network $g_{\theta}$ predicts the corresponding normalized flow $V_{t}$ with 3D coordinates $\textbf{x}=(x,y,t)$. For interpolation usage, $V_{t}$ is then transferred into bilateral flows $F_{t\rightarrow 0},F_{t\rightarrow 1}$ through denormalization.
@@
Motion latent. To achieve generalizable implicit modeling of motion dynamics, we learn implicit neural representations for motions with conditions on instance-specific motion latent inputs $L_{t}(x,y)$, which provides instance-specific motion priors at spatiotemporal coordinates $\textbf{x}=(x,y,t)$ for target motion $V_{t}(x,y)$. Specifically, we introduce a Motion Encoder to extract motion features ${K}_{i}$ from normalized flows $V_{i}$. To avoid potential temporal inconsistencies in normalized flows, we derive time-dependent motion features $K_{t}$ for given timestep $t$ from both motion features ${K}_{0}$ and ${K}_{1}$. To realize this, we apply forward warping $\overrightarrow{\omega}$ process to map every location of motion features ${K}_{i}$ to the target timestep $t$:
$K_{i\rightarrow t}=\overrightarrow{\omega}(K_{i},F_{i\rightarrow t},Z_{i}),$ (6)
$K_{i\rightarrow t}(x,y)=\sum_{a,b}Z_{i}(a,b)\cdot K_{i}(a,b),$ (7)
with acquiring warped features $K_{i\rightarrow t}(x,y)$ at the location $(x,y)$ includes pairs of spatial coordinates $(a,b)$ under the constraint of $F_{i\rightarrow t}(a,b)=(x,y)-(a,b)$. Here, $K_{i\rightarrow t}$ are warped features with timesteps $i\in\{0,1\}$ that integrate the correspondences from source (input) timesteps to the target timestep, $Z_{i}$ represents splatting weights, and $F_{i\rightarrow t}$ denotes scaled bidirectional flows at timestep $t$. We compute scaled bidirectional flows $F_{i\rightarrow t}$ as follows:
@@
Implicit motion prediction. To realize INRs for generalizable motion modeling, we devise an adaptive coordinate-based network $g_{\theta}$ for implicit and continuous motion encoding, which maps spatiotemporal coordinates $\textbf{x}=(x,y,t)$ along with the corresponding motion latent code $L_{t}$ to predicted normalized flows $\hat{V_{t}}$:
@@ 4.1 Motion modeling
Baselines. We compare GIMM with other motion modeling approaches, including Linear, Forward Warp, and End-to-End. Specifically, Linear replaces our GIMM module with the linear approximation strategy [21]. Similarly, Forward Warp substitutes GIMM with a forward warping strategy [33]. Meanwhile, the End-to-End denotes the strategy that directly predicts motion at arbitrary timesteps through an end-to-end fitting. In this end-to-end setting, we select the current state-of-the-art EMA-VFI [52] as the representative method. For fair comparisons, we employ RAFT [47] as the flow estimator for different motion modeling methods.
Evaluation settings. We measure the modeled motion quality on both VTF and VSF benchmarks, by calculating PSNR on normalized flows and End-Point-Error (EPE) on the unscaled flows. For the downstream interpolation task, we calculate PSNR on the ‘hard’ split of the SNU-FILM-arb benchmark (SNU-FILM-arb-Hard). Details about SNU-FILM-arb are presented in Section 4.3.
Table 1: Comparisons of different motion modeling methods. We assess the modeled motion on Vimeo-Triplet-Flow (VTF) and Vimeo-Septuplet-Flow (VSF) by employing PSNR and EPE metrics. Additionally, we demonstrate their impact on the interpolation task by presenting the PSNR values of their interpolation results on the ‘hard’ split of the SNU-FILM-arb dataset.
@@
Results. We provide quantitative comparisons of our GIMM and baselines in Table 1. We first report the results of motion quality on VTF and VSF. We find that GIMM can continuously model motions in videos. GIMM predicts the best flows on both the VTF benchmark (37.56dB PSNR/ 0.34 EPE) and the VSF benchmark (30.45dB PSNR/ 2.68 EPE) that involves unseen timesteps during training. Moreover, our study highlights the importance of motion priors in both motion modeling and the downstream interpolation task. The Linear, Forward Warp, and our GIMM methods that leverage motion priors from RAFT [47] can provide better motion and interpolation performances than the End-to-End across all benchmarks. Notably, GIMM benefits the interpolation task the most and achieves the highest PSNR of 32.62 dB on SNU-FILM-arb-Hard.
Visualizations. Besides the quantitative evaluation of motion modeling, we also qualitatively evaluate them on SNU-FILM-arb, as shown in Figure 4. We observe the capacity of GIMM for accurate motion modeling at arbitrary timesteps. Take the Somersault in Figure 4 for example, the ankle flow and the corresponding interpolation generated by GIMM have the best consistency with the ground truth, pointing to the same direction. Moreover, GIMM is capable of reducing ambiguities of moving objects. We can observe clearer silhouettes with fewer blurs in both the ‘dog-leg’ and ‘dog-head’ cases in Figure 4.
4.2 Ablation Studies
We investigate the effectiveness of model designs in GIMM in Table 2. The experiments are conducted on the Vimeo-Triplet-Flow (VTF) and Vimeo-Septuplet-Flow (VSF).

### v4

@@ Abstract
Motion modeling is critical in flow-based Video Frame Interpolation (VFI). Existing paradigms either consider linear combinations of bidirectional flows or directly predict bilateral flows for given timestamps without exploring favorable motion priors, thus lacking the capability of effectively modeling spatiotemporal dynamics in real-world videos. To address this limitation, in this study, we introduce Generalizable Implicit Motion Modeling (GIMM), a novel and effective approach to motion modeling for VFI. Specifically, to enable GIMM as an effective motion modeling paradigm, we design a motion encoding pipeline to model spatiotemporal motion latent from bidirectional flows extracted from pre-trained flow estimators, effectively representing input-specific motion priors. Then, we implicitly predict arbitrary-timestep optical flows within two adjacent input frames via an adaptive coordinate-based neural network, with spatiotemporal coordinates and motion latent as inputs. Our GIMM can be easily integrated with existing flow-based VFI works by supplying accurately modeled motion. We show that GIMM performs better than the current state of the art on standard VFI benchmarks.
@@ 1 Introduction
Figure 1: Schematic of motion modeling paradigms in video frame interpolation. (a) A naïve linear combination of bidirectional flows $F_{0\rightarrow 1},F_{1\rightarrow 0}$ (i.e., flows between input frames) may lead to ambiguous and coarse motion estimation due to strong overlapped and linear assumptions. (b) A time-condition-based modeling approach may predict suboptimal bilateral flows $F_{t\rightarrow 0},F_{t\rightarrow 1}$ (i.e., flows between estimated and input frames), capturing spatiotemporal changes for moving objects ineffectively. (c) Our generalizable implicit motion modeling properly represents spatiotemporal dynamics across videos and predict better bilateral flows via an adaptive coordinate-based neural network.
@@
Accurately modeling flow from two distant frames is challenging due to the complexities of capturing subtle and dynamic movements caused by varying motion speeds, object occlusions, and lighting conditions. A commonly used approach in the literature combines bidirectional flows derived from input frames [22, 40, 45, 53, 39] (Figure 1(a)). This method assumes overlapped and linear motion for flow estimation, which does not accurately reflect real-world dynamics. Several studies [20, 29, 55, 42] directly predict bilateral flows based on intermediate discrete timestamps (Figure 1(b)). These approaches model the correlation between frame motions and timestamps without leveraging motion priors within input frames; they thus fall short of capturing complex spatial-temporal changes and handling occluded regions and unexpected deformations. Moreover, this discrete-time-based modeling paradigm is ineffective for arbitrary-time interpolation.
@@
In this study, we explore a more effective and generalizable implicit approach to motion modeling for VFI. Inspired by the success of implicit neural representations [48] in encoding complex high-dimensional data such as 2D images [5], 3D scenes [34], and videos [3], we propose to implicitly model optical flows between two adjacent video frames using coordinate-based neural networks. This implicit paradigm can directly take arbitrary spatiotemporal coordinates as inputs and effectively decode the desired space-time outputs, making it a promising framework for learning highly dynamic optical flows in real-world videos. However, leveraging implicit neural networks for effective motion modeling poses challenges. First, standard implicit neural networks typically perform per-instance modeling, optimizing model parameters for a single specific input. This limitation restricts their applicability across different input video frames. Therefore, we need to develop a more adaptive implicit model capable of capturing motions in any given video. Second, efficiently integrating spatiotemporal information within implicit neural networks is complex, especially when dealing with the intricate motions occurring between two video frames. This necessitates designing appropriate implicit neural architectures that can accurately predict and represent both the spatial and temporal dynamics in videos.
@@
To this end, we propose a novel generalizable implicit flow encoding for motion modeling in VFI, called Generalizable Implicit Motion Modeling (GIMM) (Figure 1(c)). Our method only assumes the availability of bidirectional flows ($F_{0\rightarrow 1},F_{1\rightarrow 0}$ ) of two input frames obtained from a pre-trained optical flow estimator (e.g., RAFT [50], FlowFormer [19]). The input flows can be noisy as this prior will be refined to estimate the bilateral flows ($F_{t\rightarrow 0},F_{t\rightarrow 1}$) between arbitrary intermediate timestamps. Our method is unique in that it can represent input-specific motion priors effectively. For instance, it can accurately capture the complex dynamics of a somersault. This is achieved by introducing a motion encoding pipeline to extract spatiotemporal motion latent from the bidirectional flows. Our method can further enable frame interpolation at arbitrary timestamps, thanks to the adaptive coordinate-based neural network that takes spatiotemporal coordinates and motion latent as inputs. This capability allows our method to generate frames at various temporal granularities, providing flexibility and precision in video frame interpolation.
@@
Our contributions are summarized as follows: We present an effective motion modeling paradigm for video frame interpolation characterized by a novel generalizable implicit motion modeling framework. Our GIMM is capable of accurately predicting optical flow for arbitrary timesteps between two adjacent video frames at any resolution, allowing seamless integration with existing flow-based VFI methods. We demonstrate the advantages of our GIMM in motion modeling for arbitrary-timestep VFI tasks, achieving state-of-the-art performances on various benchmarks.
@@ 2 Related Work
Implicit neural representations. Implicit Neural Representations (INRs) have been shown effective in modeling complex, high-dimensional data for various applications, including video compression [3], novel view synthesis [34, 30], and image super-resolution [5]. Typically, INRs learn a continuous mapping from a set of coordinates to a specific signal using a coordinate-based neural network to encode data implicitly. The flexible and expressive modeling capabilities of INRs motivate us to explore their use to encode intricate motions and capture subtle, dynamic movements of objects in real-world videos. A recent work related to ours is IFE [13]. While both approaches consider implicit flow encoding, our method significantly differs from IFE. Unlike IFE, which focuses on per-scene encoding—where each coordinate-based network is parameterized by a specific video instance—we aim to develop a generalizable motion modeling approach that can be applied across different videos.
@@ 3 Method
Given a pair of adjacent video frames $I_{0},I_{1}\in\mathbb{R}^{H\times W\times 3}$ with timesteps $\{0,1\}$, a general flow-based video frame interpolation process is defined as follows:
@@ 3.1 Generalizable Implicit Motion Modeling
Figure 2: Our GIMM first transforms initial bidirectional flows $F_{0\rightarrow 1},F_{1\rightarrow 0}$ as normalized flows $V_{0},V_{1}$. The motion encoder extracts motion features $K_{0},K_{1}$ from $V_{0},V_{1}$, respectively. $K_{0},K_{1}$ are then forward warped at a given timestep $t$ using bidirectional flows to obtain the warped features $K_{t\rightarrow 0},K_{t\rightarrow 1}$. We pass the warped and initial motion features into a Latent Refiner that outputs motion latent $L_{t}$, representing motion information at $t$. Conditioned on $L_{t}(x,y)$, the coordinate-based network $g_{\theta}$ predicts the corresponding normalized flow $V_{t}$ with 3D coordinates $\textbf{x}=(x,y,t)$. For interpolation usage, $V_{t}$ is then transferred into bilateral flows $F_{t\rightarrow 0},F_{t\rightarrow 1}$ through denormalization.
@@
Motion latent. To achieve generalizable implicit modeling of motion dynamics, we learn implicit neural representations for motions with conditions on instance-specific motion latent inputs $L_{t}(x,y)$, which provides instance-specific motion priors at spatiotemporal coordinates $\textbf{x}=(x,y,t)$ for target motion $V_{t}(x,y)$. Specifically, we introduce a Motion Encoder to extract motion features ${K}_{i}$ from normalized flows $V_{i}$. To maintain spatiotemporal consistencies of modeled motion, we derive time-dependent motion features $K_{t}$ for given timestep $t$ from both motion features ${K}_{0}$ and ${K}_{1}$. To realize this, we apply forward warping $\overrightarrow{\omega}$ process to map every location of motion features ${K}_{i}$ to the target timestep $t$:
$K_{i\rightarrow t}=\overrightarrow{\omega}(K_{i},F_{i\rightarrow t},Z_{i}).$ (6)
Here, $K_{i\rightarrow t}$ are warped features with timesteps $i\in\{0,1\}$ that integrate the correspondences from source (input) timesteps to the target timestep, $Z_{i}$ represents splatting weights, and $F_{i\rightarrow t}$ denotes scaled bidirectional flows at timestep $t$. We compute scaled bidirectional flows $F_{i\rightarrow t}$ as follows:
@@
Implicit motion prediction. To realize INRs for generalizable motion modeling, we devise an adaptive coordinate-based network $g_{\theta}$ for implicit and continuous motion encoding, which maps spatiotemporal coordinates $\textbf{x}\in[-1,1]^{H\times W\times 3}$ along with the corresponding $D$-dimension motion latent code $L_{t}\in\mathbb{R}^{H\times W\times D}$ to predicted normalized flows $\hat{V_{t}}$:
@@ 4.1 Motion Modeling
Baselines. We compare GIMM with other motion modeling approaches, including Linear, Forward Warp, End-to-End, and BMBC [39]. Specifically, Linear replaces our GIMM module with the linear approximation strategy [22]. Similarly, Forward Warp substitutes GIMM with a forward warping strategy [35]. Meanwhile, the End-to-End denotes the strategy that directly predicts motion at arbitrary timesteps through an end-to-end fitting. In this end-to-end setting, we select the current state-of-the-art EMA-VFI [55] as the representative method. Additionally, we make further comparisons with BMBC [39], which is supposed to model complex motion effectively through several stages of motion predictions and motion refinements. For fair comparisons, we employ RAFT [50] as the flow estimator for different motion modeling methods.
Evaluation settings. We measure the modeled motion quality on both VTF and VSF benchmarks by calculating PSNR on normalized flows and End-Point-Error (EPE) on the unscaled flows. For the downstream interpolation task, we calculate PSNR on the ‘hard’ split of the SNU-FILM-arb benchmark (SNU-FILM-arb-Hard). Details about SNU-FILM-arb are presented in Section 4.4.
Table 1: Comparisons of different motion modeling methods. We assess the modeled motion on Vimeo-Triplet-Flow (VTF) and Vimeo-Septuplet-Flow (VSF) by employing PSNR and EPE metrics. Additionally, we demonstrate their impact on the interpolation task by presenting the PSNR values of their interpolation results on the ‘hard’ split of the SNU-FILM-arb dataset. We denote the calculated PSNR values as PSNRf for optical flows in the motion assessment and PSNRi for interpolated images in the interpolation task.
@@
Results. We provide quantitative comparisons of our GIMM and baselines in Table 1. We first report the results of motion quality on VTF and VSF. We find that GIMM can continuously model motions in videos. GIMM predicts the best flows on both the VTF benchmark (37.56dB PSNR/ 0.34 EPE) and the VSF benchmark (30.45dB PSNR/ 2.68 EPE) that involves unseen timesteps during training. Moreover, our study highlights the importance of motion priors in both motion modeling and the downstream interpolation task. The Linear, Forward Warp, and our GIMM methods that leverage motion priors from RAFT [50] can provide better motion and interpolation performances than the End-to-End and BMBC [39] across all benchmarks. Notably, GIMM benefits the interpolation task the most and achieves the highest PSNR of 32.62 dB on SNU-FILM-arb-Hard. While motion modeling plays a critical role in interpolation, other modules also impact interpolation outcomes. The End-to-End method has a narrower performance gap with other advanced motion modeling approaches on the interpolation benchmark than on motion modeling benchmarks, likely due to its more complex and extensive feature extraction and frame synthesis modules.
Visualizations. Besides the quantitative evaluation of motion modeling, we also qualitatively evaluate them on SNU-FILM-arb-Hard, as shown in Figure 4. Since SNU-FILM-arb-Hard requires 8X interpolation for input frames of large motion, classical methods like BMBC [39] may struggle to model the intermediate motion. In contrast, we observe the capacity of GIMM for accurate motion modeling at arbitrary timesteps. Take the Somersault in Figure 4 for example, the ankle flow and the corresponding interpolation generated by GIMM have the best consistency with the ground truth, pointing to the same direction. Moreover, GIMM is capable of reducing ambiguities when interpolating moving objects. We can observe clearer silhouettes with fewer blurs in both the ‘dog-leg’ and ‘dog-head’ cases in Figure 4.
Table 2: Quantitative comparisons of different motion model choices.
Model Variants Vimeo-Triplet-Flow (VTF) Vimeo-Septuplet-Flow (VSF) Param.(M)
[…]
Meta-learning (INR) 30.19 0.88 24.50 6.80 43.92
GIMM (U-Net) 36.96 0.39 29.96 2.90 4.27
GIMM (INR) 37.56 0.34 30.45 2.68 0.25
4.2 Ablation Study on Motion Model Choices
As our core contribution, we present an effective motion modeling paradigm for video frame interpolation, characterized by a novel generalizable implicit motion modeling framework (GIMM). In this section, we further investigate the necessity of implicit neural representations (INRs) and approaches to generalizable INRs. The experiments are conducted on the motion modeling benchmarks, i.e., Vimeo-Triplet-Flow (VTF) and Vimeo-Septuplet-Flow (VSF).
Necessity of INRs. A straightforward replacement of INRs in our GIMM is a timestep-conditioned U-Net as in diffusion literature [43]. We substitute the INRs with a U-Net of a comparable number of channels to ensure fair comparisons. In Table 2, replacing INR with U-Net results in worse performance, especially on the 6X motion modeling benchmark VSF. This demonstrates the strong continuous modeling ability of INR. Besides, GIMM with INR has a much lighter architecture, improving efficiency. Therefore, it is necessary and proper to use INR for our motion modeling.
Approaches to generalizable INRs. To realize generalizable implicit modeling, an alternative way is to modulate the weights of INRs with meta-learners [6, 25]. Following  [25], we employ Transformers [51] as the meta-learner. Compared to GIMM, this meta-learning approach performs significantly worse across all benchmarks while introducing a model with over 170X more parameters.
4.3 Ablation Study on GIMM
In this section, we investigate the effectiveness of model designs in GIMM in Table 3. The experiments are conducted on the VTF and VSF benchmarks.